OpenStudy (anonymous):

Find the solutions to the systems 9. y=x^2+5x+6 y=4x+12 A. (2,20) and (-3,0) B. (2,20) and (-3,-18) C. (-2,-20) and (-3,-18) D. no solutions 10. y=x^2-3x-1 y=8x-1 A. (0,-1) and (11,388) B. (0,-1) and (11,87) C. (-1,0) and (87,11) D. no solutions

OpenStudy (anonymous):

@ganeshie8 @ParthKohli help please(:

Parth (parthkohli):

\[y =x^2 + 5x+6\]\[y = 4x + 12\]Basically, since \(y\) is equal to both of these things, we have\[x^2 + 5x + 6 = 4x + 12\]Arrange all the terms on one side and you'll have a quadratic equation with you.

OpenStudy (ankitshaw):



OpenStudy (johnweldon1993):

y=x^2+5x+6 y=4x+12 Set them equal to each other x^2 + 5x + 6 = 4x + 12 combine like terms x^2 + x - 6 = 0 Factor out (x + 3)(x - 2) = 0 x would be -3 and 2

Parth (parthkohli):

You'll have to similarly work with the other question.

OpenStudy (johnweldon1993):

Then of course...you would plug those 2 x-values you have received into your equations again to find out what the y-coordinate will be

OpenStudy (johnweldon1993):

Similarly with number 10 y=x^2-3x-1 y=8x-1 Set them equal to each other x^2 - 3x - 1 = 8x - 1 And continue on from there combining like terms and factoring :)

OpenStudy (anonymous):

thanks everyone <3 @xxraider365 @johnweldon1993 @ParthKohli

OpenStudy (anonymous):

oh and @ankitshaw :)